# Open Your Ears to Take a Look: A State-of-the-Art Report on the Integration of Sonification and Visualization

### v1

@@ 3.1 Thematic Corpus Overview
the visualization idiom, such as scatter plot or line chart
Identity channels
@@
name of the visual channel such as color hue or shape, that is used to communicate the identity of an item (i.e., "What" something is)
Magnitude channels
@@
name of the visual channel such as position or length that is used to communicate the magnitude of an attribute (i.e., "How much" something is)
Sonification Design
Technique
the sonification technique such as parameter mapping or earcons
Identity channels
@@
the auditory channel such as timbre that is used to communicate the identity of an item (i.e., "What" something is)
Magnitude channel
@@
the auditory channel such as pitch or duration that is used to communicate the magnitude of an attribute (i.e., "How much" something is)
Reading Level
  50     29
Whole
designs that display all datapoints
  13     29
Group
designs that display a group of datapoints
  4      17
Single
designs that display single datapoints
Search Level
  2      2
Lookup
@@
the user knows the location and the target of the search
  9      6
Browse
the user knows the location but not the target of the search
  7      7
Locate
the user doesn’t know the location but the target of the search
  40     42
Explore
the user doesn’t know the location or the target of the search
  11     8
None
none of the above
Dataset Type
  29     40
Table
@@
data constructed from items and attributes (spreadsheets)
  5      3
Network
@@
data constructed from items (nodes), links and potentially attributes
  9      7
Field
data constructed from grids (positions) and attributes
  20     9
Geometry
data constructed from items and positions
Level of Measurement
  22     13
Nominal
@@
data that builds categories (such as different fruits)
  8      8
Ordinal
@@
data that builds ordered categories (such as t-shirt sizes)
  27     31
Interval
@@
data that has equal intervals, such as the time on the clock
  27     34
Ratio
@@
data that has equal intervals and a meaningful zero point, such as length or weight
Level of Redundancy
  17
Redundant
@@
a design mapping all displayed information to both senses redundantly
  14
Complementary
@@
a design mapping part of the information exclusively to the visualization and another part of the information exclusively to the sonification
  28
Mixed
a design mapping some information redundantly, some information complementary
Evaluation System
  12
User Performance
@@
evaluations collecting metrics such as error rates or task completion times
  11
User Experience
@@
evaluations collecting user feedback, typically done in a usability test
  2
Algorithmic Performance
@@
evaluations doing measurements without users such as rendering speed
  17
Qualitative Result Inspection
@@
evaluations providing subjective arguments on the quality of the result
  20
None
no evaluation done in the paper
Target Platform
  43
Desktop Computer Display
conventional screen on desktop computer
  4
XR
extended reality settings such as virtual or augmented reality glasses
  10
Physical Environments
environments that foster collaboration such as a CAVE
  5
Touch Display
@@
interactive screen that users can interact with via touch
User Interaction
  37
Yes
@@
designs that require user interaction other than pressing play
  20
No
designs that require no user interaction
Users
  25
Domain Experts
domain experts that are not visualization or sonification researchers
  7
Researchers
visualization or sonification researchers
  31
General Public
the interested general public
Goals
  8
Education
idioms designed for education
  35
Data Analysis
idioms designed for data analysis (incl. data exploration)
  3
Research
idioms designed for visualization and sonification research
  17
Public Engagement
idioms designed for public engagement
Demo
  29
Yes
@@
the paper links to a demo such as an interactive website, a video, or an audio recording
  23
No
the paper doesn’t link to a demo
  5
Yes, but not online
the paper provides a link that is not online anymore
@@
Table 2: Descriptions and distributions of all used classification tags. For the categories with two separate tags under the ”Num." column, the left and right boxes represent the visualization and sonification distributions, respectively. For the rows regarding idiom design, we report the number of different entries in the database.
@@ 3.2 Purpose
In this section, we will discuss two different purposes an audiovisual idiom can be designed for, inspired by the taxonomy described in [Mun15]: exploration and presentation. The purpose should be read as general design goal of a tool with broad perspective. Therefore, the term exploration also covers what is widely referred to as data analysis.
@@ 3.2.1 Exploration
A wide range of audiovisual idioms in our database support users with the exploration of data, see Table 2. We refer to the purpose of exploration in cases where a user intends to acquire new knowledge from their data by using an audiovisual display idiom.
Exploratory Data Analysis is the core aim of model-based sonification [Her02]. The Mode Explorer is an interactive audiovisual display idiom, combining a scatterplot visualization and model-based sonification to explore high-dimensional data [YH18]. To explore the data a user can “scratch” a scatter plot of dimensionality-reduced data with an interactive pen. The scratching introduces a virtual particle to the high-dimensional space which will follow a gradient descent towards the nearest mode in the data. The “kinetic energy” of this virtual particle is what a user will be able to hear. While the particle travels closer and closer to its final mode, the sound gradually turns more and more harmonic, finally resulting in a clear pitch when a mode is reached.
AIive combines visualization and sonification in a virtual reality environment [LLW21]. The idiom supports users in understanding the basic concepts of neural networks. Users can manipulate the weights of the nodes of a virtually displayed 3D neural network by dragging the nodes in 3D space. In real-time, the sonification displays the loss and accuracy of the neural network, therefore enabling a user to explore different constellations of node weights. Users are also able to add or delete nodes, hence experimenting with the complexity of the neural network. While the authors do not provide a user study in their paper, it is most plausible the exploratory character of the design can support people in understanding the basics of neural networks.
In a recent publication, Lemmon et al. presented an audiovisual map idiom that seeks to tackle some of the sociotechnical challenges associated with epidemiological mapping [$*{}^{*}$23]. Using their idiom, users can interactively explore Suffolk County’s experience with Covid-19. To do so, the black population and associated case numbers are displayed on the left audio channel, while the white population and associated case numbers are displayed on the right channel via the pitch of sine and triangle waves. A correlation between ethnicity and Covid-19 case numbers and their dependency on different regions becomes clearly apparent when listening to the sonification while brushing the map using the computer mouse.
@@ 3.2.2 Presentation
We refer to the purpose of presentation when a user intends to present prior knowledge to others, or in cases where a user is presented with information that is new to them but not to the designer.
@@
Recent works delve into environmental concerns and the promotion of awareness. One such noteworthy contribution is presented in [HK22]. This publication centers around a multimodal museum installation designed to foster public engagement with wildfire forecasts in specifically chosen California and South Korea regions. The project incorporates interactive data visualization, sonification, and 3D-printed sculptures. Combining these elements conveys wildfire data, creating a comprehensive and immersive experience. The installation allows audiences to explore the representation freely through contactless interaction. Without user engagement, the wildfire data representation seamlessly loops across the screen. Notably showcased at the ARKO Art Center in Seoul, this project successfully captivated the attention of 20 users. The outcomes revealed a heightened level of interest and comprehension regarding the impact of wildfires through the effective utilization of multimodal interaction. The project emphasizes observations and prompts audience reflection without demanding immediate action, leveraging new media to enhance public climate change comprehension.
@@ 3.2.3 Exploration and Presentation
Some of the designs we identified were designed both to explore and to present data. Inclusion in this category requires the designer to put similar weight on the exploration and presentation of the data. To present a few examples, Paté et al. [$*{}^{*}$22] demonstrate their audiovisual display idiom with three seismic data sets, where the sonification methods are adapted to the specific properties of each data set. A multi-scale audification method is presented where different speed factors are used depending on the size of each dataset, and different sound designs are used to highlight specific properties of the data set. This classification can also indicate that the user can explore several presentations that have been created from the same dataset and make comparisons between them. Huppenkothen et al. [HPDW23] created an audiovisual version of the Hertzsprung-Russell Diagram where the user can listen and compare auditory representations of each type of star that is included in the diagram (see Figure 7(c)). It is also possible to filter the diagram according to certain criteria, which in turn filters which sonifications are able to be listened to. Elmquist et al. [EEBR21] created complementing sonifications of the planets in the solar system which were integrated with a visualization, where a user can explore the properties of the planets and make comparisons between them. The user is able to listen to all of the sonifications for each planet at the same time, or enable specific sonifications for each planets to compare specific properties. Du et al. [$*{}^{*}$18] conducted a study investigating the enhanced visualization of basketball player movement data during a game. The visualization is designed to convey the offensive and defensive dynamics of the team. The interface primarily enables users to specify a particular time range of interest, providing a more comprehensive view of information within that timeframe. Notably, sonification is integrated exclusively during the exploration phase.
@@ 3.3 Audiovisual Idiom Design
Designers of audiovisual idioms have a sheer endless number of possibilities to combine their visualizations and sonifications. They can choose from a vast number of established visualization idioms and sonification techniques. Having chosen two designs to integrate with each other, they are free to choose the visual and auditory channels they want to map their data to. During the classification of the surveyed literature, we used six different categories to capture the state-of-the-art of audiovisual idiom design. The categories are the visualization idiom, the sonification technique, the visual identity- and magnitude channels, as well as the auditory identity- and magnitude channels. The constructs of identity- and magnitude channels are used to distinguish between encodings that communicate “what” something is and “how much of” something there is [Mun15, $*{}^{*}$23]. Typical visual identity channels are the shape of a visual mark or its color hue. Typical visual magnitude channels are the position or the length of a visual mark. A typical auditory identity channel is the timbre of a sound, often generated using different musical instruments. Typical auditory magnitude channels are pitch and loudness [DB13, CLR22]. In the following, we will make use of these descriptions to shed light on some of the designs in our database.
@@
Figure 2: The SoniScope audiovisual display idiom [$*{}^{*}$22a]. Users can interactively select a region in a scatterplot to sonify an additional data attribute for the respective items.
@@
SoniScope [$*{}^{*}$22a] is an interactive audiovisual display idiom that combines a scatterplot visualization with an interactive parameter mapping sonification (compare Figure 2). Inspired by the technique of auscultation, the SoniScope acts as a stethoscope for data, providing an auditory lens to “listen into” one’s data. The exploratory character of the idiom is most apparent when the sonification displays non-visual data dimensions. To interact with the idiom, a user can brush the data using a visual lens. Clicking into the scatterplot will then trigger the sonification of a non-visual attribute of the selected data items. Regarding the used magnitude channels, we see a common combination of visual position and the pitch of the sounds, also employed in publications such as [BF12, DLVDCG22, FBC12, FN18, KLTW17].
@@
Bearman [Bea11] studied the possibilities to display the uncertainty in future climate projections for the UK. To do so, the proposed audiovisual idiom shows a map of the UK with a heat map overlay displaying the projected temperature values. This visualization is combined with the sonification technique of interactive parameter mapping. People in a user study were asked to hover over the heat map with a mouse, triggering the sonification of the respective region. A higher pitch of a trumpet sound communicated a higher uncertainty for that region. Hence, for their design, Bearman used the spatial position as a visual identity channel and color as visual magnitude channel. This sonification did not use an identity channel (only one sound could be heard) but used pitch as magnitude channel.
@@
Figure 3: Examples of frameworks for sonification in combination with visualization. (a) The Sonification Workstation by Phillips and Cabrera [PC19] designed for general sonification tasks. (b) The WebAudioXML Sonification Toolkit (WAST) user interface by Lindetorp and Falkenberg [LF21] also aimed for general sonification tasks.(c) The Sonifigrapher virtual synthesizer interface by Riber [Rib19] for sonifying light curves. All screenshots CC BY 4.0.
@@
In their paper, Winters et al. [WW15] describe the design of the visualization and the sonification of the 2011 Tohoku earthquake in Japan. The design combines four line charts of four seismographic recording stations in Japan with an audification of the same data streams. The essence of the audification is that the low frequency recordings of the earthquake are played back in a faster tempo, making them audible to the human ear. This very direct mapping between the physical (non-audible) phenomenon and the audification to the audible range results in a rich auditory impression that becomes informative in an ecologically valid manner. The paper mostly discusses the popularity of this sonification on YouTube, currently with around 90k views, explicitly mentioning the relevance of combining the sonification with visualization for its success in public outreach. The visualization uses position as both its identity and its magnitude channel. The sonification employs the spatial stereo position of the four channels to distinguish between them, hence as their identity channel. The magnitude channels are a mix of pitch and timbre, resulting from the direct mapping between the physical phenomenon and the sound.
@@
Disscussion: From the data, we see that parameter mapping is by far the most used sonification technique for the design of audiovisual idioms – 53 out of 57 papers in the database employ some sort of parameter mapping sonification. Only a seven of them combine parameter mapping sonification with other techniques such as audification [MMU16, HPDW23], earcons [HAR16, NSC16], or auditory icons [BBV16, $*{}^{*}$15, PBM15]. The technique of audification is part of the database three times on its own [$*{}^{*}$22, HCTP14, WW15], and in combination with earcons once in a second design of a paper [$*{}^{*}$22]. Yang and Hermann are the only authors employing model-based sonification with their Mode Explorer design [YH17]. The most prominent auditory identity channel in our literature corpus is timbre (used 21 times) and the most used auditory magnitude channel is pitch (47 times), also used in combination with other channels. These findings are just in line with other meta analyses of the field of sonification [CLR22, DB13]. The visualizations in our corpus employ idioms such as line charts (10 times), scatter plots (8 times), maps (7 times), as well as networks, bar charts, slicing, heatmaps, and several other idioms. These publications use the identity channels color hue (27 times), position (11 times), as well as shape (11 times). The employed visual magnitude channels are position (33 times), color hue (15 times), size (9 times), and other channels such as opacity, tilt, or animation.
@@
Tagging the surveyed literature made us realize that with techniques such as audification or model-based sonification using the concept of the “channel” is not trivial. With these techniques, the designer of a sonification, to some extent, loses control over the sonic outcome of a sonification. This contradicts the definition of a visual channel as “a way to control the appearance of marks” [Mun15]. Instead, thinking of the “channel” as “the quality of a mark that transports information,” also allows for the description of audifications and model-based sonifications. While these techniques typically result in highly complex sound sequences, it will often be qualities such as pitch or timbre that are informative to the listener. In general, we see a quite diverse field of different audiovisual idiom designs in our corpus.

### v2

@@ 11 Both designs used for both purposes above
Visualization Design Idiom the visualization idiom, such as scatter plot or line chart
@@
Identity channels the name of the visual channel, such as color hue, or shape, that is used to communicate the identity of an item (i.e., "What" something is)
@@
Magnitude channels the name of the visual channel such as position or length that is used to communicate the magnitude of an attribute (i.e., "How much" something is)
Sonification Design Technique the sonification technique such as parameter mapping sonification or earcons
@@
Identity channels the auditory channel such as timbre that is used to communicate the identity of an item (i.e., "What" something is)
@@
Magnitude channel the auditory channel such as pitch or duration that is used to communicate the magnitude of an attribute (i.e., "How much" something is)
Reading Level   50      29 Whole designs that display all datapoints
  13      29 Group designs that display a group of datapoints
  4       17 Single designs that display single datapoints
@@
Search Level   2       2 Lookup the user knows the location and the target of the search (see Figure 4)
  9       6 Browse the user knows the location but not the target of the search
  7       7 Locate the user doesn’t know the location but the target of the search
  40      42 Explore the user doesn’t know the location or the target of the search
  11      8 None none of the above
@@
Dataset Type   29      40 Table data constructed from items and attributes (spreadsheets; see Figure 5)
@@
  5       3 Network data constructed from items (nodes), links and potentially attributes
  9       7 Field data constructed from grids (positions) and attributes
  20      9 Geometry data constructed from items and positions
@@
Level of Measurement   22      13 Nominal data that builds categories (such as different fruits)
@@
  8       8 Ordinal data that builds ordered categories (such as t-shirt sizes)
@@
  27      31 Interval data that has equal intervals, such as the time on the clock
@@
  27      34 Ratio data that has equal intervals and a meaningful zero point, such as length or weight
@@
Level of Redundancy    17 Redundant a design mapping all displayed information to both senses redundantly (see Figure 6)
@@
   14 Complementary a design mapping part of the information exclusively to the visualization and another part of the information exclusively to the sonification
   28 Mixed a design mapping some information redundantly, some information complementary
@@
Evaluation System    12 User Performance evaluations collecting metrics such as error rates or task completion times
@@
   11 User Experience evaluations collecting user feedback, typically done in a usability test
@@
   2 Algorithmic Performance evaluations doing measurements without users such as rendering speed
@@
   17 Qualitative Result Inspection evaluations providing subjective arguments on the quality of the result
   20 None no evaluation
Target Platform    43 Desktop Computer Display conventional screen on desktop computer
   4 XR extended reality settings such as virtual or augmented reality glasses
   10 Physical Environments environments that foster collaboration such as a CAVE
@@
   5 Touch Display interactive screen that users can interact with via touch
@@
User Interaction    37 Yes designs that require user interaction other than pressing “play”
   20 No designs that require no user interaction
Users    25 Domain Experts domain experts that are not visualization or sonification researchers
   7 Researchers visualization or sonification researchers
   31 General Public the interested general public
Goals    8 Education idioms designed for education
   35 Data Analysis idioms designed for data analysis (incl. data exploration)
   3 Research idioms designed for visualization and sonification research
   17 Public Engagement idioms designed for public engagement
@@
Demo    29 Yes the paper links to a demo such as an interactive website, a video, or an audio recording
   23 No the paper doesn’t link to a demo
   5 Yes, but not online the paper provides a link that is not online anymore
@@
Table 2: Descriptions and distributions of all used classification tags. For the categories with two separate tags under the ”Num." column, the left and right boxes represent the visualization and sonification distributions, respectively.
@@ 3.2 Purpose
In this section, we will discuss two different purposes an audiovisual idiom can be designed for, inspired by the taxonomy described in [Mun15]: exploration and presentation. The purpose should be read as the general design goal of a tool with a broad perspective. Therefore, the term exploration also covers what is widely referred to as data analysis.
@@ 3.2.1 Exploration
A wide range of audiovisual idioms in our database support users with the exploration of data (see Table 2). We refer to the purpose of exploration in cases where a user intends to acquire new knowledge from their data by using an audiovisual display idiom.
Exploratory data analysis is the core aim of model-based sonification [Her02]. The Mode Explorer is an interactive audiovisual display idiom, combining a scatterplot visualization and model-based sonification to explore high-dimensional data [YH18]. To explore the data, a user can “scratch” a scatter plot of dimensionality-reduced data with an interactive pen. The scratching introduces a virtual particle to the high-dimensional space, which will follow a gradient descent to the nearest mode in the data. A user will be able to hear the “kinetic energy” of this virtual particle. While the particle travels closer and closer to its final mode, the sound gradually turns more and more harmonic, finally resulting in a clear pitch when a mode is reached.
AIive combines visualization and sonification in a virtual reality environment [LLW21]. The idiom helps users understand the basic concepts of neural networks. Users can manipulate the weights of the nodes of a virtually displayed 3D neural network by dragging the nodes in 3D space. In real-time, the sonification displays the loss and accuracy of the neural network, therefore enabling a user to explore different constellations of node weights. Users are also able to add or delete nodes, hence experimenting with the complexity of the neural network. While the authors do not provide a user study in their paper, it is most plausible the exploratory character of the design can support people in understanding the basics of neural networks.
In a recent publication, Lemmon et al. presented an audiovisual map idiom that seeks to tackle some of the sociotechnical challenges associated with epidemiological mapping [LSB23]. Using their idiom, users can interactively explore Suffolk County’s experience with Covid-19. The black population and associated case numbers are displayed on the left audio channel, while the white population and associated case numbers are displayed on the right channel via the pitch of sine and triangle waves. A correlation between black and white population numbers and COVID-19 case numbers and their dependency on different regions becomes clearly apparent when listening to the sonification while brushing the map using the computer mouse.
@@ 3.2.2 Presentation
We refer to the purpose of presentation when a user intends to present prior knowledge to others or in cases where a user is presented with information that is new to them but not to the designer.
@@
Recent works delve into environmental concerns and the promotion of awareness. One such noteworthy contribution is presented in [HK22]. This publication centers around a multimodal museum installation designed to foster public engagement with wildfire forecasts in specifically chosen California and South Korea regions. The project incorporates interactive data visualization, sonification, and 3D-printed sculptures. Combining these elements conveys wildfire data, creating a comprehensive and immersive experience. The installation allows audiences to explore the representation freely through contactless interaction. Without user engagement, the wildfire data representation seamlessly loops across the screen. Showcased at the ARKO Art Center in Seoul, this project successfully captivated the attention of 20 users. The outcomes revealed a heightened level of interest and comprehension regarding the impact of wildfires through the effective utilization of multimodal interaction. The project emphasizes observations and prompts audience reflection, leveraging new media to enhance public climate change comprehension.
@@ 3.2.3 Exploration and Presentation
Some of the designs we identified were designed both to explore and to present data. Inclusion in this category requires the designer to put similar weight on the exploration and presentation of the data.
To present a few examples, Paté et al. [PFH22] demonstrate their audiovisual display idiom with three seismic data sets, where the sonification methods are adapted to the specific properties of each data set. A multi-scale audification method is presented where different speed factors are used depending on the size of each dataset, and different sound designs are used to highlight specific properties of the data. Huppenkothen et al. [HPDW23] created an audiovisual version of the Hertzsprung-Russell Diagram where the user can listen and compare auditory representations of each type of star that is included in the diagram (see Figure 7(c)). It is also possible to filter the diagram according to certain criteria, which, in turn, filters which sonifications can be listened to. Elmquist et al. [EEBR21] created complementing sonifications of the planets in the solar system, which were integrated with a visualization, where a user can explore the properties of the planets and make comparisons between them. The user can listen to all of the sonifications for each planet at the same time or enable specific sonifications for each planet to compare specific properties. Du et al. [DCM18] conducted a study investigating the enhanced visualization of basketball player movement data during a game. The visualization is designed to convey the offensive and defensive dynamics of a team. The interface primarily enables users to specify a particular time range of interest, providing a more comprehensive view of information within that timeframe. Notably, sonification is integrated exclusively during the exploration phase.
@@ 3.3 Audiovisual Idiom Design
Designers of audiovisual idioms have countless possibilities to combine their visualizations and sonifications. They can choose from a vast number of established visualization idioms and sonification techniques. Having chosen two designs to integrate with each other, they are free to choose the visual and auditory channels they want to map their data to. During the classification of the surveyed literature, we used six different categories to capture the state-of-the-art of audiovisual idiom design. The categories are the visualization idiom, the sonification technique, the visual identity- and magnitude channels, as well as the auditory identity- and magnitude channels. The constructs of identity- and magnitude channels are used to distinguish between encodings that communicate “what” something is and “how much of” something there is [Mun15, ERI23]. Typical visual identity channels are the shape of a visual mark or its color hue. Typical visual magnitude channels are the position or the length of a visual mark. A typical auditory identity channel is the timbre of a sound, often generated using different musical instruments. Typical auditory magnitude channels are pitch and loudness [DB13, CLR22]. In the following, we will make use of these descriptions to shed light on some of the designs in our database.
@@
Figure 2: The SoniScope audiovisual display idiom [ERI22]. Users can interactively select a region in a scatterplot to sonify an additional data attribute for the respective items.
@@
SoniScope [ERI22] is an interactive audiovisual display idiom that combines a scatterplot visualization with an interactive parameter mapping sonification (compare Figure 2). Inspired by the technique of auscultation, the SoniScope acts as a stethoscope for data, providing an auditory lens to “listen into” one’s data. The exploratory character of the idiom is most apparent when the sonification displays non-visual data dimensions. To interact with the idiom, a user can brush the data using a visual lens. Clicking into the scatterplot will then trigger the sonification of a non-visual attribute of the selected data items. Regarding the used magnitude channels, we see a common combination of visual position and the pitch of the sounds, also employed in publications such as [BF12, DLVDCG22, FBC12, FN18, KLTW17].
@@
Bearman [Bea11] studied the possibilities of displaying the uncertainty in future climate projections for the UK. The proposed audiovisual idiom shows a map of the UK with a heat map overlay displaying the projected temperature values. This visualization is combined with the sonification technique of interactive parameter mapping. People in a user study were asked to hover over the heat map with a mouse, triggering the sonification of the respective region. A higher pitch of a trumpet sound communicated a higher uncertainty for that region. Hence, for their design, Bearman used the spatial position as a visual identity channel and color as a visual magnitude channel. This sonification did not use an identity channel (only one sound could be heard) but used pitch as a magnitude channel.
@@
Figure 3: Examples of frameworks for sonification in combination with visualization. (a) The Sonification Workstation by Phillips and Cabrera [PC19] is designed for general sonification tasks. (b) The WebAudioXML Sonification Toolkit (WAST) user interface by Lindetorp and Falkenberg [LF21] also aimed for general sonification tasks. (c) The Sonifigrapher virtual synthesizer interface by Riber [Rib19] for sonifying light curves. All screenshots CC BY 4.0.
@@
In their paper, Winters et al. [WW15] describe the design of the visualization and the sonification of the 2011 Tohoku earthquake in Japan. The design combines four line charts of four seismographic recording stations in Japan with an audification of the same data streams. The essence of the audification is that the low-frequency recordings of the earthquake are played back at a faster tempo, making them audible to the human ear. This very direct mapping between the physical (non-audible) phenomenon and the audification to the audible range results in a rich auditory impression that becomes informative in an ecologically valid manner. The paper mostly discusses the popularity of this sonification on YouTube, currently with around 90k views, explicitly mentioning the relevance of combining the sonification with visualization for its success in public outreach. The visualization uses position as both its identity and its magnitude channel. The sonification employs the spatial stereo position of the four channels to distinguish between them, hence as their identity channel. The magnitude channels are a mix of pitch and timbre, resulting from the direct mapping between the physical phenomenon and the sound.
@@
Discussion: From the data, we see that parameter mapping is by far the most used sonification technique for the design of audiovisual idioms – 53 out of 57 papers in the database employ some sort of parameter mapping sonification. Only seven of them combine parameter mapping sonification with other techniques such as audification [MMU16, HPDW23], earcons [HAR16, NSC16], or auditory icons [BBV16, RFK15, PBM15]. The technique of audification is part of the database three times on its own [PFH22, HCTP14, WW15], and once in combination with earcons in a second design of a paper [PFH22]. Yang and Hermann are the only authors employing model-based sonification with their Mode Explorer design [YH17]. The most prominent auditory identity channel in our literature corpus is timbre (used 21 times), and the most used auditory magnitude channel is pitch (47 times), also used in combination with other channels. These findings are just in line with other meta-analyses of the field of sonification [CLR22, DB13]. The visualizations in our corpus employ idioms such as line charts (10 times), scatter plots (8 times), and maps (7 times), as well as networks, bar charts, heatmaps, and several other idioms. These publications use the identity channels color hue (27 times), position (11 times), as well as shape (11 times). The employed visual magnitude channels are position (33 times), color hue (15 times), size (9 times), and other channels such as opacity, tilt, or animation.
@@
Tagging the surveyed literature made us realize that using the concept of the “channel” for techniques such as audification or model-based sonification is not trivial. With these techniques, the designer of a sonification, to some extent, loses control over its sonic outcome. This contradicts the definition of a visual channel as “a way to control the appearance of marks” [Mun15]. Instead, thinking of the “channel” as “the quality of a mark that transports information,” also allows for the description of audifications and model-based sonifications. While these techniques typically result in highly complex sound sequences, it will often be qualities such as pitch or timbre that are informative to the listener. In general, we see a quite diverse field of different audiovisual idiom designs in our corpus.